Two blocks of masses 2Kg & 1Kg are tied to a string which passess over a light pulley.Initially the masses are held at rest on the same horizontal line.Find the distance travelled by centre of mass in 2 sec__________

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear sindhuja p

acceleration of masses is given by a=(m2-m1)g/(m2+m1)

                                                   =g/3

1 kg moves in upword direction  =1/2 a t2

                                              =1/2 *g/3 *4

                                               =2g/3

2kg moves in downward direction =1/2 a t2

                                                =2g/3

Now center of mass ={1*2g/3  +2*(-2g/3)}/(1+2)

                            =-2g/9

 so  center of mass 2g/9 m in downward direction


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin